# Semantic Mapping in Indoor Embodied AI – A Comprehensive Survey and Future Directions

### v1

@@ 7.2 Dense yet efficient maps
Following our discussion on various map structures, we find it crucial that the semantic map be able to capture dense visual cues to allow for complex spatial reasoning among objects. For example, beyond addressing straightforward queries like ‘Where is the table?’, the semantic maps can be assessed on more intricate spatial reasoning queries such as ‘Can you retrieve my phone from my desk beside my laptop?’. To perform such reasoning, the semantic map needs to capture fine-grained detail about the spatial arrangement of the objects (phone and laptop on the table) in the scene. While at one end of the spectrum topological maps are too sparse failing to capture the dense semantics of a scene, at the other end the point-cloud based maps are too dense capturing redundant empty space information. Although spatial top-down 2D maps exist somewhere in the middle, they fail to capture the semantic information in the third dimension (height). This is crucial where the navigation is in 3D space, for example in drones. Hence there is still a need of a dense-enough map to capture the 3D space in its entirety at the same time intelligently ignoring empty space information. Additionally, a dense map will consume more memory and will be difficult to update. So a dense map representation which is still scalable, memory and computation efficient is a research direction worthwhile to pursue.
@@ 8 Conclusion
In this survey, we explore various approaches to semantic map building in the Embodied AI literature, focusing on indoor environments. Existing works primarily employ spatial maps to capture geometric layouts or topological maps to model landmark-based relationships. While many robotics studies have explored hybrid maps that combine spatial and landmark information, this approach remains under-explored in embodied AI research. It presents a promising avenue for future work, particularly in leveraging such maps to enhance performance on complex spatial reasoning tasks. We also discuss how dense point-cloud maps, created by associating semantic information with point clouds or triangle meshes, offer potential for embodied AI tasks but remain under-explored. While promising for spatial reasoning, their high memory demands, computational inefficiency, and the presence of mostly empty 3D space in indoor environments pose significant challenges. With recent advances in large foundation models, the focus has shifted towards building open-vocabulary, queryable map representations that are task-agnostic. Despite these advances, learning to map an indoor 3D scene in simulation rely on unrealistic assumptions such as perfect localization and noiseless sensors, leading to a significant sim-to-real gap. Moreover, these approaches assume static environments, limiting their applicability in dynamic outdoor settings like autonomous driving.
Building on the insights from this survey, we highlight the limitations in existing semantic map-building methods and explore promising directions for future research. We aim for this discussion to not only summarize the current state of the field but also inspire the research community to advance it further.
@@ Acknowledgements
The authors were supported by Canada CIFAR AI Chair grant and NSERC Discovery Grant. We also thank Dr. Manolis Savva, Dr. Yasutaka Furukawa, Tommaso Campari, Austin T. Wang and Xingguang Yan for their valuable feedback on our paper.

### v2

@@ 8.2 Dense yet efficient maps
Following our discussion on various map structures, we find it crucial that the semantic map be able to capture dense visual cues to allow for complex spatial reasoning among objects. For example, beyond addressing straightforward queries like ‘Where is the table?’, the semantic maps can be assessed on more intricate spatial reasoning queries such as ‘Can you retrieve my phone from my desk beside my laptop?’. To perform such reasoning, the semantic map needs to capture fine-grained detail about the spatial arrangement of the objects (phone and laptop on the table) in the scene. While at one end of the spectrum topological maps are too sparse failing to capture the dense semantics of a scene, at the other end the dense geometric maps are too dense capturing redundant empty space information. Although spatial top-down 2D maps exist somewhere in the middle, they fail to capture the semantic information in the third dimension (height). This is crucial where the navigation is in 3D space, for example in drones. Hence there is still a need of a dense-enough map to capture the 3D space in its entirety at the same time intelligently ignoring empty space information. Additionally, a dense map will consume more memory and will be difficult to update. So a dense map representation which is still scalable, memory and computation efficient is a research direction worthwhile to pursue.
@@ 8.5 Devising evaluation metrics
8.6 Summary
This section outlines promising future directions, as derived from the current challenges in semantic mapping in Sec. 7. As the field moves toward creating flexible, general-purpose, and queryable maps, researchers face open problems in balancing semantic richness with efficiency, ensuring scalability and robustness in dynamic environments, and developing better evaluation metrics that go beyond task performance to assess map quality directly. These provide a multitude of future research directions, which will help advance the field.
@@ 9 Conclusion
In this survey, we review a wide range of semantic map-building approaches and categorize them based on their underlying map structure, such as spatial grids, topological graphs, dense geometric and hybrid representations, as well as the nature of semantic encoding, such as explicit or implicit. This perspective is timely in the light of recent advances in foundation models and the increasing need for general-purpose, multi-modal, open-vocabulary and queryable map representations. This survey helps identify key challenges in current semantic mapping paradigms and present promising directions for future research in semantic mapping.
Existing works primarily employ spatial maps to capture geometric layouts or topological maps to model landmark-based relationships. While many robotics studies have explored hybrid maps that combine spatial and landmark information, this approach remains under-explored in embodied AI research. It presents a promising avenue for future work, particularly in leveraging such maps to enhance performance on complex spatial reasoning tasks. We also discuss how dense geometric maps, created by associating semantic information with point clouds, triangle meshes or surfels, offer potential for embodied AI tasks but remain under-explored. While promising for spatial reasoning, their high memory demands, computational inefficiency, and the presence of mostly empty 3D space in indoor environments pose significant challenges. Moreover, most current mapping approaches assume static environments, limiting their applicability in dynamic settings, for example in scenes where furniture is rearranged or humans move around.
We hope that the insights presented in this survey serve to guide and inspire further advancements in the field by the research community.
@@ Acknowledgements
The authors were supported by Canada CIFAR AI Chair grant and NSERC Discovery Grant. We also thank Manolis Savva, Yasutaka Furukawa, Tommaso Campari, Austin T. Wang, Xingguang Yan, Bernadette Bucher, Duy Ta and Sachini Herath for their valuable feedback on our paper.